Express the solution to the following inequality using interval notation.
x^2 + 7x - 18 ≥ 0

Answer by tommyt3rd(5050)

x^2+7x-18=
(x+9)(x-2)

and since a=1 (a>0) we know that it opens upwards.
this implies that it is negative on (-9,2)
therefore our solution is...
(-oo,-9]U[2,oo)